OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Which of the following sums would be under the radical symbol to find the distance between the points (7, -1) and (-8, -9)? A. 1^2 + 10^2 B. 1^2 + 8^2 C. 15^2 + 8^2

OpenStudy (anonymous):

the distance formula is \[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\], where x1, x2, y1 and y2 are coordinates of points.

OpenStudy (anonymous):

So do I plug the numbers into the formula?

OpenStudy (anonymous):

\[7-(-8)=15\] \[-1-(-9)=-1+9=8\]

OpenStudy (anonymous):

yes just plug in

OpenStudy (anonymous):

under the radical symbol there will be \[15^2+8^2\], so the answer is C

OpenStudy (anonymous):

Oh, thank you so much! :)

OpenStudy (anonymous):

np :)